MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2016/08/17
OBSERVER: Alex Kris Paul
CONDITION: clear with high cloud
COMMENT:
Compressor gas pressure: 1.8 MPa
Expected Observation Time: START 19:01 END 06:41


Go to the Seeing Log

Go to the Observation Log

TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB55876-559032826116040
LMC13282-132909466060
SMC-000
TRANSIT-000
ALERT-000
GRB-000
TOTAL---3730722100 (06h08m20s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
910262426269252566
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
924569955528


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
231112888223
Field13Field14Field15Field16
2300


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000


TRANSIT FIELD STATISTICS

Field1Field2Field3
000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
19:01observingclear, thin cloud around
20:00observingclear, thin cloud around
21:00observingclear, thin cloud around
22:00observingcloudy, slightly clear
23:00observingcloudy, slightly clear
00:00observingclear, thin cloud around
01:00observingclear, thin cloud around
02:00observingheavy cloud
03:00observingcloudy, slightly clear
04:00observingcloudy, slightly clear
05:00observingcloudy, slightly clear
06:00waitingcloudy, slightly clear
06:41waitingcloudy, slightly clear


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day